OpenStudy (lonababy):

please help!!! Solve, using the substitution method. m + n = 5 m – n = 3 A. There is no solution. B. The solution is (2, –1) C. The solution is (4, 1) D. There are an infinite number of solutions.

OpenStudy (blurbendy):

You can use the first equation, m + n = 5 to solve for m, then use that answer and substitute it into the second equation. This is just one of many approaches.

OpenStudy (lonababy):

thank you!

OpenStudy (blurbendy):

Do you understand?

OpenStudy (lonababy):

yea kinda

OpenStudy (blurbendy):

okay, what does m equal?

OpenStudy (blurbendy):

in the first equation

OpenStudy (lonababy):

5?? idk

OpenStudy (blurbendy):

you just subtract n from both sides to get m by itself.

OpenStudy (lonababy):

ok is it like switching it around and then instead of keeping the + u put -?

OpenStudy (blurbendy):

you need to get m by itself on the left side first, but there is an + n, so you have have to subtract it from both sides: m + n = 5 -n -n m = 5 - n The +n and -n on the right side cancel each other out, so you are left with 5 - n on the right side. Make sense?

OpenStudy (lonababy):

ohh ok I think I get it now

OpenStudy (blurbendy):

good

OpenStudy (lonababy):

I have no idea

OpenStudy (blurbendy):

i did that wrong

OpenStudy (blurbendy):

m - n = 3 (5 -n) - n = 3 5 - n - n = 3 5 -2n = 3 -2n = -2 n = 1 now substitute this value back into the first equation m + 1 = 5 So, what does m equal?